Moon1619:

Suppose that there are two types of tickets to a show: advance and same-day. The combined cost of one advance ticket and one same-day ticket is $75. For one performance, 15 advanced tickets and 20 same-day tickets were sold. The total amount paid for the tickets was $1300. What was the price of each kind of ticket?

AZ:

Hi @moon1619

AZ:

Do you know how to come up with the equations to solve this question?

Moon1619:

@AZ kind of I struggle with math is it 75x+15=1300?

AZ:

That's alright! And unfortunately it isn't We need to come up with TWO separate equations and then solve it

AZ:

So we have two different type of tickets an advance ticket and a same-day ticket do you want to choose a letter (or variable) for it? we can stick with x or y, unless you want to use something else like A for advance ticket and S for same-day ticket

CripQUEZZ:

So X + Y = $75 Once you find the (X) or the advanced and the (Y) or the same day you multiply (X) by (15) and (Y) by (20)

Moon1619:

@AZ I use X and Y

Moon1619:

@AZ @cripQUEZZ would the first one be 15x+20y=1300

CripQUEZZ:

Well first we gotta find wat x and y equal

DADDYROXY:

obv

AZ:

That's correct!

AZ:

And like @CripQUEZZ already said, the other equation would be x + y = 75

CripQUEZZ:

Yeah

AZ:

so x is our advanced tickets y is our same-day tickets

AZ:

So do you know how to solve a system of equations? There's three methods 1. Elimination method 2. Substitution method 3. Graphing

AZ:

Are you familiar with any of these?

Moon1619:

@AZ yes

AZ:

Which one would you like to use? If you already know how to do the substitution and elimination method and you're just trying to solve it faster, the graphing method is the fastest way and we can use https://www.desmos.com/calculator

CripQUEZZ:

So your two equations are 15x + 20y = 1300 And X + Y = 75

CripQUEZZ:

Do the first equation then you can get he second one easily

AZ:

The point where the two lines intersect would be the solution to the two equatons the x-value giving you the price for advanced tickets and the y-value giving you the price for same-day tickets

Moon1619:

@AZ so if I graph 15 and 20 whatever the intercept is is the answer?

AZ:

You graph the two lines x+y=75 15x+20y=1300 Where the two lines intersects gives you the answer. The intercepts are different than where they INTERSECT (different words)

AZ:

But only use this graphing method if you're comfortable with solving the system of equations by either the substitution or elimination method It's important to know how to do it by yourself first :)

Moon1619:

@AZ oh ok I don’t understand how to graph it but to solve this would I do 15x+20y=1300 subtract 20 from 1300 then divided by 15

AZ:

(also you don't have to @ me in every message since I get a notification each time you respond)

AZ:

we can't subtract 20 because there's a 'y' with it we have to solve the two equations together

CripQUEZZ:

No not really youll divide both numbers

AZ:

so let's do it step by step first we have x + y = 75 can you solve this for x?

Moon1619:

(Oh ok) yes but we have to rearrange it so X would equal 75 I think

AZ:

no no no think of x as apples and y as bananas you can't just throw away the bananas and say that there are 75 apples

Moon1619:

So then we can’t solve for X so we move to 15x+20y=1300

AZ:

x + y = 75 we have two DIFFERENT letters here to get 'x' all by itself, we have to subtract y on both sides and we get x + y - y = 75 - y banana minus banana = no banana so on the left hand side of the equal sign, the `y` gets cancelled out and on the right hand side, we keep the y and so we're left with x = 75 - y

AZ:

does that make sense?

Moon1619:

Yes

AZ:

so now we can use the substitution method and we'll plug that into the other equation let me add some colors for you

AZ:

\(\color{orange}{x} = \color{orange}{75-y}\) \( 15\color{orange}{x}+20y=1300\) so we can replace the `x` in the second equation with `75-y` from the first equation because we said that x = 75 - y so 75-y is the same thing as x so we can replace it this is called the substitution method

AZ:

still following along or do you have any questions?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ \(\color{orange}{x} = \color{orange}{75-y}\) \( 15\color{orange}{x}+20y=1300\) so we can replace the `x` in the second equation with `75-y` from the first equation because we said that x = 75 - y so 75-y is the same thing as x so we can replace it this is called the substitution method \(\color{#0cbb34}{\text{End of Quote}}\) very very good job !!!

Moon1619:

I kind of understand that but still confused because now the X isn’t there

AZ:

so we can use ANY letter we can use A or B or X or Y and these are just variables just imagine that y as your NEW `x` and you just solve for `y` as you would solve for `x`

AZ:

Does that help?

Moon1619:

So would it be 15 (75-y)+20y=1300

AZ:

You got it! now just solve for `y`

Moon1619:

Y= 35 and x=40

AZ:

Great job!!

Moon1619:

The only thing is I don’t know what x represents and what y represents

Moon1619:

Would x be the advance tickets?

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @AZ so x is our advanced tickets y is our same-day tickets \(\color{#0cbb34}{\text{End of Quote}}\)

Moon1619:

Ok thank you so much

AZ:

Yes! Because as you noticed in our equation 15x+20y=1300 we're multiply 15 with x and 15 was how many advanced tickets they bought x is going to be `the price of each advanced ticket` and similarly y will be the `price of each same-day ticket`

AZ:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Moon1619 Ok thank you so much \(\color{#0cbb34}{\text{End of Quote}}\) You're most welcome!